A 1.803 g sample of gypsum, a hydrated salt of calcium sulfate, CaSO4 is heated at a temperature greater than 170 degree Celsius in a crucible until a constant mass is reached. The mass of anhydrous CaSO4 salt is 1.426g. Calculate the percent by mass of water in the hydrated calcium sulfate salt.
1
Expert's answer
2018-09-27T05:47:17-0400
m(H2O) = 1.803 - 1.426 = 0.377 g
the percent by mass of water in the hydrated calcium sulfate salt = 0.377*100%/1.803 = 20.91%